# Kuan-Cheng Chen QuEST Imperial College London London, United Kingdom

### v2

@@ Abstract
This paper investigates the application of Quantum Support Vector Machines (QSVMs) with an emphasis on the computational advancements enabled by NVIDIA’s cuQuantum SDK, especially leveraging the cuTensorNet library. We present a simulation workflow that substantially diminishes computational overhead, as evidenced by our experiments, from exponential to quadratic cost. While state vector simulations become infeasible for qubit counts over 50, our evaluation demonstrates that cuTensorNet speeds up simulations to be completed within seconds on the NVIDIA A100 GPU, even for qubit counts approaching 784. By employing multi-GPU processing with Message Passing Interface (MPI), we document a marked decrease in computation times, effectively demonstrating the strong linear speedup of our approach for increasing data sizes. This enables QSVMs to operate efficiently on High-Performance Computing (HPC) systems, thereby opening a new window for researchers to explore complex quantum algorithms that have not yet been investigated. In accuracy assessments, our QSVM achieves up to 95% on challenging classifications within the MNIST dataset for training sets larger than 100 instances, surpassing the capabilities of classical SVMs. These advancements position cuTensorNet within the cuQuantum SDK as a pivotal tool for scaling quantum machine learning simulations and potentially signpost the seamless integration of such computational strategies as pivotal within the Quantum-HPC ecosystem.
Index Terms:
Quantum Machine Learning, Quantum Simulation, Tensor Network, cuQuantum SDK, cuTensorNet
I Introduction
@@
Enter the realm of quantum computing, a burgeoning field offering profound computational speedups over classical approaches for certain problem types. Among these, Quantum Support Vector Machines (QSVMs) emerge as a promising quantum-enhanced technique for machine learning[6, 7], capable of drastically reducing the computational resources required for SVMs. Leveraging quantum algorithms, QSVMs achieve exponential speedups in both training and classification tasks by performing calculations in parallel and employing quantum-specific optimizations[6, 8, 9, 10].
@@
However, in the current Noisy Intermediate-Scale Quantum (NISQ) era[11], the practical utility of quantum computers is significantly constrained by their availability and imperfect technological state. Challenges such as the fidelity of qubits, the error rates of two-qubit gates, and the limited number of available qubits present substantial hurdles[12, 13, 14]. Despite the advent of several methodologies aimed at enhancing qubit fidelity—such as Quantum Error Mitigation (QEM)[15, 16] and Dynamical Decoupling (DD)[17, 18]—these limitations persist, impeding the realization of quantum advantage on quantum computing platforms in the current NISQ era[19, 20]. Consequently, the design and validation of quantum-inspired algorithms, or hybrid classical-quantum algorithms, are predominantly conducted through high-performance classical simulations[10, 21]. Furthermore, quantum simulators have shown considerable success in the near-term verification of quantum algorithms on small qubit systems[22].
@@
Within the scope of our research, we have engineered an advanced tensor-network simulation framework, purpose-built to expedite the development of QSVMs through the integration of the cuTensorNet library underlying cuQuantum SDK[23]. This library is meticulously optimized for NVIDIA GPUs and can facilitate QSVM algorithms, requiring noiseless simulations for quantum kernel estimation as depicted in Fig. 1. A pre-computation mechanism is embedded within this workflow, allowing for the reuse of quantum kernel values in the QSVM’s complex learning stages, thereby bolstering the efficacy of both the training and classification phases.
Our simulation of the QSVM algorithm, devised for parallel execution using the Message Passing Interface (MPI), harnesses the formidable power of GPU acceleration, equipping our QSVM simulator to efficiently manage vast data sets with only a modest increment in memory requirements. Its flexibility ensures its utility across various quantum machine learning paradigms. Performance benchmarks of our simulator demonstrate that it achieves significant speedups, often exceeding tens of times, which are exponentially better than those achieved by state-of-the-art methods using Qiskit on CPUs[24]. Furthermore, it boasts scalability to multi-GPUs—a testament to the proficiency achieved by the NVIDIA Quantum Team—thereby affirming our simulator’s role as a potent and scalable asset in the quantum machine learning arena and the Quantum-HPC ecosystem[23, 21].
@@
This approach’s adeptness at handling up to 784 qubits for QSVMs allows for a comprehensive scaling analysis, shedding light on the quantum acceleration achieved in comparison to traditional classical solvers. These strides in QSVM development signal a major progression towards practical deployment, charting a path for the application of quantum-enhanced methodologies to complex, real-world data classification challenges within the Quantum-HPC Ecosystem[10, 21, 25, 26, 27]. This paves the way for breakthroughs in quantum computing’s applicability and marks a significant contribution to the quantum information sciences.
II Background
In the research work by Rebentrost et. al.[6], a quantum algorithm is introduced that revolutionizes the computational efficiency of SVMs for big data classification. This QSVM leverages quantum mechanics to achieve logarithmic complexity in terms of both the dimension of the vectors $N$ (qubit number) and the number of training examples $M$ (data size).
Traditionally, SVM is formulated to identify a hyperplane that maximizes the margin between two classes, described in the primal form as:
@@ II Background
$y_{j}(\mathbf{w}\cdot\mathbf{x}_{j}+b)\geq 1,\;\forall j,$ (2)
where $\mathbf{w}$ is the normal to the hyperplane, $b$ is the bias, $\mathbf{x}_{j}$ are the feature vectors, and $y_{j}$ are the class labels.
The quantum version utilizes the kernel method, mapping the feature space to a higher-dimensional space to allow for nonlinear classification. The kernel matrix $K$, with elements $K_{jk}=k(\mathbf{x}_{j},\mathbf{x}_{k})$, represents inner products in this space. The quantum algorithm efficiently inverts this kernel matrix, a task that is computationally intensive in classical SVM, especially as $M$ and $N$ grow.
The quantum algorithm for matrix inversion, central to QSVM, significantly reduces the runtime complexity to $O(\log(NM))$[6]. This quantum routine for non-sparse matrix exponentiation underpins the SVM’s optimization, facilitated by the QML algorithm for linear equations, known as the HHL algorithm, which solves $A|x\rangle=|b\rangle$ for Hermitian matrix $A$.
The SVM’s dual optimization problem is given by:
$L(\alpha)=\sum_{j=1}^{M}y_{j}\alpha_{j}-\frac{1}{2}\sum_{j,k=1}^{M}\alpha_{j}K_{jk}\alpha_{k},$ (3)
with constraints:
$\sum_{j=1}^{M}\alpha_{j}y_{j}$ $=0,$ (4)
$\alpha_{j}$ $\geq 0,\;\forall j.$ (5)
The QSVM algorithm employs quantum routines to approximate the solution vector $\alpha$, defining the separating hyperplane in the feature space, thus classifying new data points with reduced computational complexity. This approach highlights the potential of quantum computing to tackle large-scale machine learning problems, characteristic of big data challenges, by utilizing quantum parallelism and efficient manipulation of high-dimensional data.
@@
III Simulating QSVM using cuQuantum SDK and cuTensorNet
III-A Introduction of cuQuantum SDK and cuTensorNet
As the fields of quantum computing and advanced numerical simulations rapidly expand, NVIDIA has introduced cuQuantumK[23], a comprehensive software development kit (SDK), to accelerate quantum circuit simulations with NVIDIA GPUs. The cuQuantum SDK consists of optimized libraries such as cuStateVec and cuTensorNet. cuStateVec is dedicated to state-vector-based simulation methods, enabling substantial acceleration and efficient memory utilization. cuTensorNet, on the other hand, deals with tensor-network-based simulation, where a quantum circuit is represented as a tensor network (Fig.2). In this formulation, the sequence of pairwise contractions plays a role in computational cost as illustrated in Fig.3. The cuTensorNet library provides advanced features for tensor network contractions, including path optimization, approximate tensor network simulation techniques, and multi-GPU multi-node execution. These functionalities facilitate simulations at an unprecedented scale with significant acceleration, empowering researchers to delve into unexplored tensor network theories and complex quantum algorithms. As the demand for more sophisticated computational methods grows within the scientific community, cuStateVec and cuTensorNet stand out for their potential to accelerate research domains across quantum physics, quantum chemistry, and quantum machine learning.
Figure 2: Quantum Circuit Formulation of cuTensorNet
@@ III-A Introduction of cuQuantum SDK and cuTensorNet
Figure 3: Consider two contraction paths: The upper path results in a higher computational cost
@@
cuTensorNet can efficiently find a contraction path with high quality quickly[23]. This capability accelerates the exploration of quantum mechanics underlying complex systems and quantum machine learning models, particularly in the processing and analysis of high-dimensional data.
Figure 4: Building blocks of cuTensorNet
Consequently, cuQuantum itself represents a comprehensive suite aimed at bolstering quantum computing simulations. As part of NVIDIA’s broader initiative to support the quantum computing ecosystem, cuQuantum provides a range of tools and libraries engineered to optimize various facets of quantum computing simulations, including the programming model CUDA Quantum[23, 28] (CUDA-Q) and frameworks such as Qiskit[29], Pennylane[30], and Cirq.[31]. The introduction of cuQuantum marks a significant stride towards bridging the theoretical potential of quantum computing with the practical realities of today’s computational resources. By offering a scalable and high-performance platform for quantum simulations, cuQuantum not only democratizes access to quantum computing research but also propels the field towards achieving real-world quantum computing applications.
III-B Pipeline of simulating QSVM algorithm
In Fig. 5(a), the depicted pipeline of a QSVM commences with the initial quantum state preparation in a canonical basis state $|0\rangle$. The number of qubits depends on input data features, which can be adjusted using principal components analysis (PCA) to evaluate QSVM with varying qubit counts. A parameterized quantum circuit (QC) follows, designed to map classical data onto a quantum feature space via tunable gates, with the gate parameters refined throughout the training process to facilitate effective data classification. This step, shown in Fig. 5(b), is counterpoised by the introduction of the adjoint of the parameterized quantum circuit (QC), which ensures the reversion of ancillary qubits to their initial state, maintaining the quantum state’s coherence for measurement. Central to the QSVM’s operation is the quantum entanglement and interference within the high-dimensional feature space, as visualized in the intertwined loops, where the data classification is a result of the constructive and destructive interference patterns that are encoded in the state’s amplitude probabilities. The measurement stage yields classical outputs indicative of the data’s class membership probabilities, signifying the culmination of the QSVM process—from quantum state preparation and transformation through to interference-based processing and eventual classical information retrieval—for enhanced machine learning efficacy. In this paper, we use a parameterized quantum circuit based on Block-Encoded State (BPS) wavefunctions[32, 33]. This enables QSVM to maintain high classification accuracy even with a greater number of qubits. Worthy of note, our quantum circuit does not decompose into small blocks; instead, we entangle each qubit by adding CNOT gates with linear topology. Such nearest-neighbor connectivity renders the QSVM circuit compatible with near-term quantum hardware.
@@ III-B Pipeline of simulating QSVM algorithm
Figure 5: (a) The QSVM pipeline showcasing the quantum circuit transformation of input data into feature space quantum states. (b) A schematic of the QSVM circuit.
III-C Complexity of Quantum circuit simulation for QSVM
@@
When executed on classical hardware such as CPUs and GPUs, the simulation of the QSVM algorithm poses significant computational challenges. Figure 6 elucidates these challenges, indicating that the complexity scales exponentially with the number of qubits as $O(2^{n})$ and quadratically with data size as $O(n^{2})$. Additionally, the complexity scales exponentially with the number of qubits $q$, which encode the feature space within the quantum circuit. This aspect underscores the inherent computational intensity of simulating quantum systems on classical infrastructure.
@@
III-D Simulating QSVM’s Kernel Matrix
In this section, we will discuss three methods for simulating a QSVM algorithm as follows. One method involves using a CPU with opt-einsum to optimize the simulation of quantum circuits on CPU devices. The other methods utilize cuStateVector and cuTensorNet for simulating quantum circuits on GPU devices, with acceleration provided by the cuQuantum SDK.
III-D1 Conventional Simulation of QSVM using CPU/opt-einsum
Optimized Einsum (opt-einsum) significantly enhances the execution efficiency of einsum-like expressions across various libraries by optimizing contraction orders and employing specialized routines such as BLAS and cuBLAS. Its compatibility with multiple backends, like NumPy, Dask, PyTorch, and TensorFlow, makes it a versatile tool for tensor operations on CPUs.
  1. 1.
  2. 2.
    Set the current operand index $i$ to $-1$.
  3. 3.
    for $i_{1},i_{2}\in\{1,\ldots,indices\}$ do
  1. (a)
    Update the operands index $i\leftarrow i+1$.
  2. (b)
    Compute amplitude $amp\leftarrow\text{opt\_einsum.contract}(exp,operands[i],path)$.
  3. (c)
    Calculate and store $kernel\_matrix[i_{1}-1][i_{2}-1]\leftarrow\sqrt{\text{amp.real}^{2}+\text{amp.imag}^{2}}$.
end for
  • 4.
    return $kernel\_matrix$
  • Algorithm 1 Get Kernel Matrix using opt-einsum
    To ensure a fair comparison between CPU and GPU performance, we utilize the opt-einsum package, which provides optimized tensor computation on CPUs similar to the cuQuantum SDK available for NVIDIA GPUs. The detailed algorithm for simulating the QSVM on CPUs, aimed at equalizing the computational environment to the extent possible, is described in Algorithm 1.
    III-D2 Simulation of QSVM using cuStateVector

### v3

# Validating Large-Scale Quantum Machine Learning: Efficient Simulation of Quantum Support Vector Machines Using Tensor Networks

@@ Abstract
We present an efficient tensor-network-based approach for simulating large-scale quantum circuits exemplified by Quantum Support Vector Machines (QSVMs). Experimentally, leveraging the cuTensorNet library on multiple GPUs, our method effectively reduces the exponential runtime growth to near-quadratic scaling with respect to the number of qubits in practical scenarios. Traditional state-vector simulations become computationally infeasible beyond approximately 50 qubits; in contrast, our simulator successfully handles QSVMs with up to 784 qubits, executing simulations within seconds on a single high-performance GPU. Furthermore, utilizing the Message Passing Interface (MPI) for multi-GPU environments, our method demonstrates strong linear scalability, effectively decreasing computation time as dataset sizes increase. We validate our framework using the MNIST and Fashion MNIST datasets, achieving successful multiclass classification and highlighting the potential of QSVMs for high-dimensional data analysis. By integrating tensor-network techniques with advanced high-performance computing resources, this work demonstrates both the feasibility and scalability of simulating large-qubit quantum machine learning models, providing a valuable validation tool within the emerging Quantum–HPC ecosystem.
  • * The first three authors contributed equally to this work.
  • December 2024
1 Introduction
@@
Enter the realm of quantum computing, a burgeoning field offering profound computational speedups over classical approaches for certain problem types. Among these, Quantum Support Vector Machines (QSVMs) emerge as a promising quantum-enhanced technique for machine learning[6, 7, 8], capable of drastically reducing the computational resources required for SVMs. Leveraging quantum algorithms, QSVMs achieve exponential speedups in both training and classification tasks by performing calculations in parallel and employing quantum-specific optimizations[6, 9, 10, 11].
@@
However, in the current Noisy Intermediate-Scale Quantum (NISQ) era[12], the practical utility of quantum computers is significantly constrained by their availability and imperfect technological state. Challenges such as the fidelity of qubits, the error rates of two-qubit gates, and the limited number of available qubits present substantial hurdles[13, 14, 15]. Despite the advent of several methodologies aimed at enhancing qubit fidelity—such as Quantum Error Mitigation[16, 17] and Dynamical Decoupling[18]—these limitations persist, impeding the realization of quantum advantage on quantum computing platforms in the current NISQ era[19, 20]. Consequently, the design and validation of quantum-inspired algorithms, or hybrid classical-quantum algorithms, are predominantly conducted through high-performance classical simulations[11, 21]. Furthermore, quantum circuit simulators have shown considerable success in the near-term verification of quantum algorithms on small qubit systems[22, 23].
@@
Within the scope of our research, we have engineered an advanced tensor-network simulation framework, purpose-built to expedite the development of QSVMs through the integration of the cuTensorNet library underlying cuQuantum SDK[24]. This library is meticulously optimized for NVIDIA GPUs and can facilitate QSVM algorithms, requiring noiseless simulations for quantum kernel estimation as depicted in Fig. 1. A pre-computation mechanism is embedded within this workflow, allowing for the reuse of an optimized tensor-network contraction path in the QSVM’s complex learning stages, thereby bolstering the efficacy of both the training and classification phases.
Our tensor-network-based simulation is designed for parallel execution using the Message Passing Interface (MPI) and leverages the substantial computational power of GPU acceleration. This combination enables our QSVM simulator to efficiently manage large datasets while only modestly increasing memory requirements, thereby avoiding out-of-memory (OOM) situations during large-scale quantum circuit simulations. Its flexibility ensures its utility across various quantum machine learning paradigms. Benchmark results show that our simulator achieves speedups often exceeding an order of magnitude compared with existing methods [25, 26], thereby underscoring its potential as a robust and scalable tool for quantum machine learning within the broader Quantum-High-Performance Computing (HPC) ecosystem [24, 21].
@@
A key feature of our simulator is its capacity to handle up to 784 qubits, enabling an extensive scaling analysis of QSVM performance and shedding light on the potential of quantum kernel methods in realistic data classification scenarios. Furthermore, this approach is flexible enough to accommodate various quantum machine learning paradigms and can be extended to multi-GPU settings for large-scale simulations. By validating our methods on real-world datasets (such as MNIST and Fashion-MNIST), we demonstrate that QSVMs can tackle complex classification tasks in Quantum-HPC environments, marking a significant step toward practical quantum-enhanced machine learning. These strides in QSVM development signal a major progression towards practical deployment, charting a path for the application of quantum-enhanced methodologies to complex, real-world data classification challenges within the Quantum-HPC Ecosystem[11, 21, 27, 28, 29].
These results highlight not only the viability of QSVM algorithms but also the value of advanced simulation tools in guiding future quantum hardware development, such as offering ground truth for benchmarking purpose. As such, this work contributes to bridging the gap between theoretical QSVM formulations and their eventual implementation on large-scale quantum devices, offering valuable insights for both algorithmic refinement and hardware optimization in the quantum information sciences.
2 Background
QSVMs represent a significant breakthrough in quantum machine learning, particularly for large-scale data classification. The pioneering work by Rebentrost et al. [6] introduced a quantum algorithm that substantially enhances the computational efficiency of traditional SVMs. By harnessing quantum-mechanical principles such as superposition and interference, QSVMs can, under certain assumptions (e.g., quantum random-access memory, qRAM), achieve near-logarithmic complexity with respect to both the dimensionality of feature vectors $N$ (qubit number) and the size of the training dataset $M$ (data size). This approach suggests a potential exponential speedup over classical methods, although practical constraints like the realization of qRAM remain a major challenge.
More recent work on quantum machine learning has shifted toward quantum kernel estimation, emphasizing the capability of entangled quantum states to embed classical data in an exponentially large Hilbert space [30]. Rather than focusing solely on matrix-inversion routines, these methods evaluate inner products of quantum states (i.e., kernel functions) that would be prohibitively expensive to compute classically. By embedding data into high-dimensional quantum feature spaces, one can construct decision boundaries that may be unreachable with purely classical algorithms. Indeed, Ref. [31] demonstrates an end-to-end quantum speedup for a suitably constructed classification problem, providing concrete evidence that quantum kernels can yield practical gains in machine learning tasks.
Classical SVMs aim to find a hyperplane that maximizes the margin between two classes, typically formulated in its primal form as:
@@ 2 Background
$y_{\text{j}}(\mathbf{w}\cdot\mathbf{x}_{\text{j}}+b)\geq 1,;\forall\text{j},$ (2)
where $\mathbf{w}$ is the normal vector to the hyperplane, $b$ is the bias, $\mathbf{x}_{j}$ are feature vectors, and $y_{j}$ are the class labels. In quantum extensions of this method, the data are mapped into a higher-dimensional Hilbert space via a quantum kernel, enabling efficient non-linear classification that would otherwise be computationally prohibitive on classical hardware.
Early QSVM formulations relied on quantum matrix-inversion routines, such as the HHL algorithm [32], to mitigate the computational bottleneck inherent to large-scale SVMs. Theoretical analyses indicated that QSVM could perform these matrix inversions with $\mathcal{O}(\log(NM))$ complexity [6], a significant improvement over classical approaches. Quantum parallelism further reduces computational overhead by allowing simultaneous calculation of many kernel matrix elements, crucial for SVM optimization.
Within the QSVM framework, data points $x_{\text{i}}$ are non-linearly transformed into quantum states $\rho(x_{\text{i}})=|\psi(x_{\text{i}})\rangle\langle\psi(x_{\text{i}})|$ within the Hilbert space. The inner product between these quantum states, crucial for constructing the kernel matrix $K(x_{\text{i}},x_{\text{j}})$, is given by:
$K(x_{\text{i}},x_{\text{j}})=\text{tr}{\rho(x_{\text{i}})\rho(x_{\text{j}})}=|\langle\psi(x_{\text{i}})|\psi(x_{\text{j}})\rangle|^{2},$ (3)
where $|\langle\psi(x_{\text{i}})|\psi(x_{\text{j}})\rangle|^{2}$ is computed using a unitary matrix $U$, defined as:
$|\langle\psi(x_{\text{i}})|\psi(x_{\text{j}})\rangle|^{2}=|\langle 0^{\otimes N}|U^{\dagger}(x_{\text{i}})U(x_{\text{j}})|0^{\otimes N}\rangle|^{2},$ (4)
with $|0^{\otimes N}\rangle$ representing the initial state with all qubits in the $|0\rangle$ state.
QSVM extends classical SVM by utilizing quantum superposition and entanglement to address large, complex datasets more efficiently. Quantum parallelism enables rapid evaluation of kernel functions across multiple data pairs, a task that is computationally expensive classically [33]. In light of these developments, modern QSVM research increasingly emphasizes quantum kernel methods, reflecting both the capabilities of near-term quantum devices and the desire to circumvent the strict requirements of fully functional qRAM. Consequently, the present work examines classical simulations of quantum kernel approaches, building on recent theoretical and experimental advances to investigate whether and how quantum-enhanced feature spaces can yield advantages in realistic classification scenarios.
@@
3 Simulating QSVM with Tensor Networks Using the cuQuantum SDK and cuTensorNet Library
3.1 Introduction of cuQuantum SDK and cuTensorNet Library
@@
Figure 2: (a) Tensor network formulation of a quantum circuit. (b) Contraction paths determine the computational and memory costs of tensor network simulations: The upper path results in higher costs.
As the fields of quantum computing and advanced numerical simulations rapidly expand, NVIDIA has introduced cuQuantum SDK[24], a comprehensive software development kit (SDK), to accelerate quantum circuit simulations with NVIDIA GPUs. It supports the programming model CUDA Quantum[24, 34] (CUDA-Q), and frameworks like Qiskit[35], Pennylane[36], and Cirq[37]. By offering a scalable and high-performance platform for quantum simulations, cuQuantum can democratize access to quantum computing research and even propel the field towards achieving real-world quantum machine learning applications.
The cuQuantum SDK consists of optimized libraries such as cuStateVec and cuTensorNet. cuStateVec is dedicated to state-vector simulation methods, providing significant acceleration and efficient memory usage, while cuTensorNet focuses on tensor-network simulations. For tensor-network methods, the quantum circuit is initially converted into a tensor-network representation (Fig.2(a)). Subsequently, pairwise contraction paths are optimized to minimize computational complexity and memory footprint, followed by the execution of the computation. As shown in Fig.2(b), the sequence of pairwise contractions plays a role in computational cost. cuTensorNet efficiently identifies high-quality contraction paths[24], accelerating quantum machine learning exploration, especially for high-dimensional data. The library offers advanced features like path optimization, approximate simulations, multi-GPU, and multi-node execution, enabling large-scale simulations and significantly advancing research into complex quantum algorithms across quantum physics, quantum chemistry, and quantum machine learning.
Figure 3: Building blocks of the cuTensorNet library.
@@
3.2 Pipeline of QSVM simulation
In Fig. 4(a), the depicted pipeline of a QSVM commences with the initial quantum state preparation in a canonical basis state $|0\rangle$. The number of qubits depends on input data features, which can be adjusted using principal components analysis (PCA) to evaluate QSVM with varying qubit counts. The QSVM circuit comprises a parameterized quantum circuit (QC) and its corresponding adjoint ($\text{QC}^{\dagger}$), which correspond to the unitary operators $U(\text{x}_{i})$ and $U^{\dagger}(\text{x}_{j})$ depicted in Fig. 4(b). The paired input data $\text{x}_{i}$ and $\text{x}_{j}$ are embedded into the left and right halves of the parameterized quantum circuit (QC and QC), as shown in Fig. 4(b). At the measurement stage, the probability amplitude of the zero state $|0\rangle$ is used to represent the similarity between $\text{x}_{i}$ and $\text{x}_{j}$ in the quantum feature space. After computing the zero state amplitude for all paired data in the quantum feature space, the quantum kernel matrix is used to train the support vector classifier. Notably, only the probability of the all-zero state needs to be computed, allowing the tensor network simulation to reduce the overall computation by contracting the subspaces of the complete tensor structure. In this paper, we use a parameterized quantum circuit based on Block-Encoded State (BPS) wavefunctions[38, 39]. This enables QSVM to maintain high classification accuracy even with a greater number of qubits. Notably, the circuit does not decompose into smaller blocks; instead, each qubit is entangled through CNOT gates arranged in a linear topology, ensuring compatibility with near-term quantum hardware.
@@ 3.2 Pipeline of QSVM simulation
Figure 4: (a) The QSVM pipeline showcases the quantum circuit transformation of input data into feature space quantum states. (b) A schematic of the QSVM circuit.
3.3 Complexity of Quantum Circuit Simulation for QSVM
@@
When executed on classical hardware such as CPUs and GPUs, the simulation of the QSVM algorithm poses significant computational challenges with state vector simulations. Figure 5 elucidates these challenges, indicating that the complexity scales exponentially with the number of qubits as $O(2^{n})$ and quadratically with data size as $O(N^{2})$. Additionally, the memory footprint of the full state vector grows exponentially with the number of qubits $q$, which map features into Hilbert space for quantum circuit simulations. This aspect underscores the inherent computational intensity of simulating quantum systems on classical infrastructure.
@@
3.4 Simulating QSVM’s Quantum Kernel Matrix
In this section, we discuss three methods for simulating a QSVM algorithm: state-vector simulation on GPU using the cuStateVec library, tensor-network simulation on CPU using the opt-einsum library, and tensor-network simulation on GPU using the cuTensorNet library. In this research work, our interest lies in comparing a state-of-the-art CPU-centric approach (opt-einsum) to a state-of-the-art GPU-centric approach (cuTensorNet). This comparison highlights how GPU acceleration can significantly impact large-scale quantum circuit simulations.
3.4.1 Simulation of QSVM with state vector
State vector simulation is widely used for simulating quantum circuit-based quantum computing because quantum circuit operations can naturally be represented using state vectors. In this method, Qiskit is used to create the QSVM quantum circuit, and the state-vector simulation is implemented on a GPU via the cuStateVec backend, as described in Algorithm 1. The advantage of using cuStateVec includes a speedup of the simulation time by leveraging GPU capabilities and enabling multi-GPU processing with MPI for distributed computing. The effectiveness of cuStateVec in enhancing quantum-circuit-simulation efficiency is evidenced in Lykov et al.’s research work using cuStateVec and the cuQuantum SDK[22].